• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name weinstein

    Log entry time 21:21:25 on October28,2002

    Entry number 89225

    This entry is a followup to: 89223

    keyword=magnet ramp problem solved

    19:00 noticed that Q1left is ramping MUCH too slowly (~8 hours to reach full current)
    19:10 sweep hall anyway, Scott will work on magnet in controlled access
    19:55 sweep complete
    Scott tried manual magnet ramping - does not ramp any faster
    still gives the error message 'Magnet mismatch'
    He's got more stuff to try but he's not very optimistic.
    20:45 problem solved by rebooting the Q1 L&R IOC

    Scot Spiegel - I Swapped the programmable ramp board in the Left Q1 Danfysik P/S. However, this did not work. Returned P/S to orig. configuration.
    I thought it might be a communications problem and decided to cycle power on
    I.O.C. HA48 located in personnal passageway. I set both Q1's amps to ZERO and let all magnet energy dump. Then reset I.O.C. HA48 (designated IOC for Q1's and Big Byte power supply). It took about 2 min. to start up and gui's to return. I reset the momentums for Q1's on Hall A General Tools control screen. The ramp rate was corrected and operated normally.